Find the change in internal energy of a system, if the system expels a net thermal energy of 2.5 × 104J from the environment, then the system does a work of 17 × 103 J on the environment. Explain your steps during solution.


1
Expert's answer
2020-08-24T13:22:19-0400

Solution.

According to the first law of thermodynamics, the amount of heat transferred to the system changes its internal energy and performs work on external bodies.

"Q = \Delta U + A"

"\Delta U = Q-A"

"\Delta U = 2.5*10^4-17*10^3 = 8kJ"

Answer:

"\Delta U = 8kJ"